# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v2

@@ G.3 Ligand Receptor Interactions
Figure 9: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for two consecutive slides.
Appendix H Additional Experiments & Ablations
H.1 Next Step Sampling for Axolotl Brain Regeneration
MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
11 $0.5$ $2.467\pm 0.107$ $2.301\pm 0.163$ $0.037\pm 0.002$ $9.532\pm 1.093$
11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
0 $0.5$ $2.447\pm 0.142$ $2.337\pm 0.216$ $0.036\pm 0.005$ $9.696\pm 1.882$
0 $0.8$ $413\pm 0.099 2.2$$293\pm 0.161$ $0.036\pm 0.002$ $9.114\pm 1.092$
$0.5$ $0.2$ $2.460\pm 0.118$ $2.342\pm 0.144$ $0.036\pm 0.003$ $9.500\pm 1.067$
$0.5$ $0.5$ $2.504\pm 0.094$ $2.309\pm 0.139$ $0.036\pm 0.003$ $9.394\pm 1.431$
$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
MOTFM $7.487\pm 0.698$ $7.449\pm 0.931$ $0.266\pm 0.010$ $81.965\pm 9.812$
11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
11 $0.8$ $695\pm 0.443 7.7$$792\pm 0.463$ $0.266\pm 0.007$ $87.179\pm 6.690$
0 $0.2$ $8.170\pm 0.663$ $8.079\pm 0.723$ $0.269\pm 0.008$ $91.572\pm 8.802$
0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
0 $0.8$ $7.382\pm 1.068$ $7.234\pm 0.852$ $0.267\pm 0.009$ $81.635\pm 14.135$
$0.5$ $0.2$ $94\pm 0.239 7.171$$71\pm 0.422$ $0.266\pm 0.001$ $78.924\pm 3.715$
11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
H.2 IVP Sampling on Axolotl Brain Regeneration
IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
11 $0.5$ $3.932\pm 0.377$ $4.356\pm 0.398$ $0.156\pm 0.025$ $31.524\pm 4.875$
11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
0 $0.5$ $3.943\pm 0.413$ $4.241\pm 0.435$ $0.150\pm 0.039$ $29.221\pm 5.713$
0 $0.8$ $3.881\pm 0.368$ $4.094\pm 0.551$ $0.139\pm 0.026$ $27.941\pm 6.676$
$0.5$ $0.2$ $4.152\pm 0.341$ $4.322\pm 0.291$ $0.166\pm 0.014$ $33.299\pm 3.629$
$0.5$ $0.5$ $4.013\pm 0.187$ $4.138\pm 0.297$ $0.153\pm 0.020$ $30.941\pm 3.685$
$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
11 $0.8$ $6.875\pm 0.643$ $6.920\pm 0.796$ $0.159\pm 0.045$ $62.838\pm 16.897$
0 $0.2$ $6.722\pm 0.905$ $6.782\pm 1.003$ $0.154\pm 0.034$ $53.996\pm 15.617$
0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
0 $0.8$ $6.504\pm 0.925$ $6.744\pm 1.336$ $0.174\pm 0.037$ $56.687\pm 18.118$
$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
H.3 Next Step Sampling for Mouse Embryo Organogenesis
Table 15: Interpolation via Next Step Sampling at holdout time 5 for the Mouse Organogenesis dataset.
11 $0.8$ $1.889\pm 0.024$ $1.888\pm 0.082$ $0.161\pm 0.002$ $11.552\pm 0.166$
0 $0.2$ $1.893\pm 0.035$ $1.912\pm 0.057$ $0.159\pm 0.001$ $11.462\pm 0.154$
0 $0.5$ $1.877\pm 0.039$ $1.933\pm 0.088$ $0.162\pm 0.002$ $11.528\pm 0.110$
$0.5$ $0.5$ $1.899\pm 0.027$ $1.899\pm 0.072$ $0.160\pm 0.002$ $11.517\pm 0.097$
CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $0.164\pm 0.001$ $11.499\pm 0.123$
11 $1.898\pm 0.029$ $1.866\pm 0.097$ $0.167\pm 0.002$ $11.795\pm 0.170$
$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $0.164\pm 0.002$ $11.639\pm 0.182$
Table 16: Extrapolation via Next Step Sampling at holdout time 8 for Mouse Organogenesis.
11 $0.5$ $1.685\pm 0.096$ $1.714\pm 0.159$ $0.089\pm 0.006$ $8.056\pm 1.033$
11 $0.8$ $1.703\pm 0.063$ $1.830\pm 0.131$ $0.095\pm 0.005$ $8.928\pm 0.723$
0 $0.2$ $1.715\pm 0.123$ $1.860\pm 0.267$ $0.094\pm 0.009$ $9.021\pm 1.740$
0 $0.5$ $1.725\pm 0.082$ $1.856\pm 0.191$ $0.093\pm 0.006$ $8.806\pm 0.749$
0 $0.8$ $1.774\pm 0.053$ $1.897\pm 0.175$ $0.094\pm 0.007$ $9.466\pm 0.957$
$0.5$ $0.2$ $1.818\pm 0.096$ $2.089\pm 0.222$ $0.084\pm 0.008$ $8.875\pm 0.976$
$0.5$ $0.5$ $1.774\pm 0.104$ $1.899\pm 0.280$ $0.093\pm 0.007$ $9.139\pm 1.437$
$0.5$ $0.8$ $1.768\pm 0.058$ $1.858\pm 0.120$ $0.101\pm 0.006$ $9.303\pm 0.634$
11 $1.890\pm 0.046$ $1.877\pm 0.103$ $0.147\pm 0.006$ $10.752\pm 0.405$
$0.5$ $1.636\pm 0.060$ $1.684\pm 0.099$ $\bm{0.081\pm 0.005}$ $7.088\pm 0.692$
H.4 IVP Sampling for Mouse Embryo Organogenesis
Extrapolating to the last holdout time point of the mouse organogenesis dataset (Chen et al., 2022), particularly under IVP-Sampling, represents the most challenging setting among all our experiments. This difficulty arises because the target time point lies entirely outside the training horizon, requiring integration from the initial samples through to the end. As a result, the velocity field has more opportunity to drift in incorrect directions, often leading to generations that deviate substantially from the true dynamics. In our experiments, this instability was evident: across 10 runs, several produced highly unstable trajectories, reflecting the sensitivity of the system to initial conditions and numerical solvers. This variability is also captured in the performance metrics reported in Table 18.
11 $0.5$ $3.137\pm 0.407$ $4.093\pm 1.187$ $0.086\pm 0.004$ $11.948\pm 1.393$
11 $0.8$ $3.392\pm 0.757$ $4.716\pm 2.079$ $0.089\pm 0.005$ $9.547\pm 0.752$
0 $0.2$ $2.953\pm 0.425$ $3.816\pm 0.973$ $0.083\pm 0.002$ $9.816\pm 0.715$
0 $0.5$ $2.938\pm 0.476$ $3.904\pm 1.120$ $0.088\pm 0.005$ $9.864\pm 0.764$
0 $0.8$ $3.101\pm 0.539$ $3.855\pm 0.946$ $0.087\pm 0.004$ $9.280\pm 0.551$
$0.5$ $0.5$ $3.090\pm 0.635$ $4.596\pm 2.357$ $0.084\pm 0.005$ $9.786\pm 1.067$
$0.5$ $0.8$ $3.200\pm 0.403$ $3.555\pm 0.637$ $0.084\pm 0.004$ $9.269\pm 0.541$
11 $5.200\pm 0.799$ $6.306\pm 1.037$ $0.123\pm 0.008$ $45.862\pm 13.765$
IVP MOTFM $110835\pm 211671$ $1021005\pm 2063905$ $0.086\pm 0.002$ $14178\pm 29475$
CTF-C 11 $0.2$ $785586\pm 1318212$ $7598321\pm 13497483$ $0.088\pm 0.002$ $98199\pm 150412$
11 $0.5$ $2691\pm 3931$ $28480\pm 36483$ $0.087\pm 0.002$ $1632\pm 2090$
11 $0.8$ $2473\pm 3349$ $19537\pm 26306$ $0.087\pm 0.003$ $517\pm 616$
0 $0.2$ $1493\pm 2497$ $14563\pm 24858$ $0.087\pm 0.001$ $800\pm 1158$
0 $0.5$ $218018\pm 471298$ $1820788\pm 3994886$ $0.086\pm 0.001$ $2170\pm 4697$
0 $0.8$ $2736\pm 34766 11$$18089\pm 310135$ $0.084\pm 0.002$ $27013\pm 60065$
$0.5$ $0.2$ $8114720\pm 16270274$ $69458305\pm 140579849$ $0.088\pm 0.002$ $901074\pm 1775139$
$0.5$ $0.5$ $2414338\pm 6009993$ $23103811\pm 56863018$ $0.086\pm 0.001$ $261335\pm 663279$
$0.5$ $0.8$ $158\pm 3023 11$$1138\pm 30025$ $0.084\pm 0.002$ $445\pm 1085$
CTF-H 0 $353428\pm 952168$ $3011396\pm 8057131$ $0.095\pm 0.004$ $22990\pm 58936$
11 $\bm{15\pm 10}$ $\bm{53\pm 53}$ $0.098\pm 0.006$ $\bm{48\pm 32}$
$0.5$ $107889\pm 275882$ $994606\pm 2772756$ $0.087\pm 0.002$ $8875\pm 24264$
H.5 Liver Regeneration
Table 19: Wasserstein distances for different model configurations
Variant $\bm{\lambda}$ $\bm{\alpha}$ $\bm{\mathcal{W}_{2}}$
EOT $34.30348\pm 1.44797$
CTF-C 11 $0.2$ $34.44455\pm 1.19306$
CTF-C 11 $0.5$ $33.95671\pm 1.64415$
CTF-C 11 $0.8$ $34.62812\pm 0.98181$
CTF-C 0 $0.2$ $34.24147\pm 1.16930$
CTF-C 0 $0.5$ $32.74147\pm 1.86351$
CTF-C 0 $0.8$ $33.71729\pm 1.23057$
CTF-C $0.5$ $0.2$ $33.56646\pm 1.04376$
CTF-C $0.5$ $0.5$ $33.84199\pm 1.71408$
CTF-C $0.5$ $0.8$ $33.04534\pm 1.64399$
CTF-H 0 $\bm{32.68215\pm 1.47185}$
CTF-H 11 $33.48050\pm 1.00149$
CTF-H $0.5$ $33.41444\pm 0.99501$
H.6 IVP Cell Type Progression over time
Figure 10: Temporal cell type predictions from ContextFlow for the major cell types in the Organogenesis dataset (Chen et al., 2022). Early progenitor populations (neural crest and mesenchyme) progressively diminish as development advances, while terminal fates (muscle, cartilage primordium, and liver) emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout. Overall, ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
@@ I.1 Ablation on $\lambda$
First, we conduct a sensitivity analysis on all three datasets for the trade-off parameter $\lambda$, which controls the relative importance of spatial smoothness (SS) and cell-cell communication (LR) priors. All the other hyperparameters are kept constant. From Tables 20-24 and their corresponding Figures 11-15, we observe that the best performance is usually achieved towards the extremities, at $\lambda=0$ or $\lambda=1$, with values near the latter dominating more often. We hypothesize that the SS prior (at $\lambda=1$ only the SS prior is considered), acting as a proxy for spatial distance between cells from different slices, always carries relevant information, encoding the structural information present in the data. On the other hand, the informativeness of cell–cell communication patterns depends on how distinct the ligand–receptor features are at a given time step compared with those of neighboring ones. When LR features remain highly similar across consecutive timesteps, they contribute little to the discriminative signal that the OT objective can leverage. Consequently, the influence of communication priors is strongly dataset- and timestep-dependent. This effect is also reflected in the observation that settings with $\lambda=0$ (where only the LR prior is used) tend to perform worse when using Next-Step sampling—where local, immediate effects dominate—than under IVP sampling, which integrates information over the entire preceding trajectory.
We therefore recommend experimenting with different values of $\lambda$ (e.g., 0, $0.8$, or $1$) depending on the specific use case and context. Due to the scalability of ContextFlow, hyperparameter exploration can be performed efficiently, allowing for a rapid assessment of the effect of $\lambda$ on model performance.
@@ I.2 Ablation on $r$
We conduct an ablation study on the Brain Regeneration dataset, examining the effect of varying the neighborhood radius $r$ used to define the boundary for computing the Spatial Smoothness Prior. We evaluate two settings: $\lambda=1$ and $\lambda=0.8$, corresponding to the use of only the spatial prior and to a setting with a modest contribution from the cell–cell communication prior, respectively. From Tables 25-28 and their corresponding Figures 16-19, we observe that the optimal neighborhood radius tends to lie toward the lower end of the tested range. Radii smaller than this optimum degrade performance by failing to capture sufficient local context, resulting in neighborhood means that are overly similar to individual cellular profiles. Conversely, increasing the radius beyond the optimal range also reduces performance, as the neighborhood begins to include cells from distinct types or spatial regions, thereby diluting the local signal. While certain deviations from this trend occur, likely reflecting underlying biological complexity, this behavior is consistent with the trade-off between spatial specificity and contextual coverage inherent to neighborhood-based priors.
For our case, we set the radius by considering the timestep with the least number of cells, dividing it by half (to account for different hemispheres), and dividing by the order of cell types present in that timestep. For the dataset considered in this study, Stage 44 had the fewest number of $1400$ cells, with approximately $10$ cell types present. We thus set the radius at $50$ in our studies.
@@ I.3 Ablation on $\epsilon$
Additionally, we conduct an ablation study on the Brain Regeneration dataset by varying the parameter $\epsilon$, which is used to weigh the entropic term in the OT objective. The ContextFlow configuration we consider here is CTF-H ($\lambda=1$), which includes only the spatial smoothness prior. As observed from Tables 29-30 and their corresponding Figures 20-21, drastically decreasing $\epsilon$ results in the OT formulation to ignore the relative entropic term containing the prior information and only to consider the transport cost resulting in higher Wasserstein values. Furthermore, in accordance with results from Theorem 3.1, increasing $\epsilon$ too much still does not drastically degrade the performance, as the prior matrix $M$ acts as a soft filter and prohibits uniform couplings.
While setting $\epsilon$ values, one must look at the Gibbs kernel used in the Sinkhorn Algorithm $\exp(-\mathbf{C}/\epsilon)$, since lower $\epsilon$ values can cause potential numerical issues. We thus set $\epsilon$ by examining the order of the cost matrix $\mathbf{C}$, and for the studies above, we set it to $100$ after considering the order of the median of all the elements in the cost matrix.
@@
0 $1.508\pm 0.047$ $1.386\pm 0.088$ $6.33\times 10^{5}\pm 1.72\times 10^{6}$ $3.95\times 10^{6}\pm 1.05\times 10^{7}$
0.2 $1.614\pm 0.081$ $1.642\pm 0.136$ $9.18\times 10^{2}\pm 1.69\times 10^{3}$ $4.29\times 10^{4}\pm 1.05\times 10^{5}$
0.5 $1.638\pm 0.069$ $1.676\pm 0.114$ $1.08\times 10^{5}\pm 2.76\times 10^{5}$ $9.95\times 10^{5}\pm 2.77\times 10^{6}$
0.8 $1.617\pm 0.042$ $80\pm 0.094 1.66×1$$6\times 10^{5}\pm 3.92\times 10^{5}$ $1.43\times 10^{6}\pm 3.43\times 10^{6}$
1 $1.906\pm 0.071$ $1.892\pm 0.092$ $1.89\times 10^{1}\pm 1.77\times 10^{1}$ $6.89\times 10^{1}\pm 7.36\times 10^{1}$
@@
Table 24: Interpolation for holdout timestep 3 with IVP Sampling on the Liver Regeneration dataset.
@@
Table 27: Extrapolation for the last holdout timestep with CTF-H at $\lambda=0.8$ on the Brain Regeneration dataset.
@@
Table 28: Interpolation for the middle holdout timestep 3 with CTF-H at $\lambda=0.8$ on the Brain Regeneration dataset.

### v3

@@ G.3 Ligand-Receptor Interactions
Figure 10: Visual translation of the bias of NPTX2–NPTXR LR pattern in cell type coupling.
G.4 Cell Type Predictions from ContextFlow
Figure 11: Temporal predictions from ContextFlow under IVP sampling for major cell types on the Mouse Organogenesis dataset. Early progenitor populations (neural crest and mesenchyme) progressively diminish as development advances, while terminal fates (muscle, cartilage primordium, and liver) emerge at later stages. Major lineages such as the brain, heart, and connective tissue remain continuous throughout. Overall, ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
Appendix H Statistical Significance Testing
We evaluate the statistical significance of ContextFlow’s performance using a paired two-sided $t$-test over repeated runs. For each metric, we compute paired differences between the best (bolded) setting from the result tables from Section 5 for CTF and the corresponding MOTFM baseline, let $\{d_{i}\}_{i=1}^{n}$ be these differences; the test statistic is given by $t=\bar{d}/(s_{d}/\sqrt{n})$, where $\bar{d}$ and $s_{d}$ are the sample mean and standard deviation of the differences. The $p$-value measures the likelihood of observing such a deviation in metric values by chance, assuming there is no true difference between the methods; the lower the $p$-value, the better. Table 10 and Table 11 report the p-value on the Brain Regeneration and Mouse Organogenesis datasets. For Liver regeneration, the p-value is computed as $2.74\mathrm{e}{-02}$.
Table 10: P-values for interpolation and extrapolation on Brain Regeneration from Tables 1 and 2.
Sampling Interpolation Extrapolation
Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step $1.53\mathrm{e}{-02}$ $3.62\mathrm{e}{-03}$ $5.64\mathrm{e}{-04}$ $3.57\mathrm{e}{-04}$ $1.07\mathrm{e}{-02}$ $2.58\mathrm{e}{-02}$ $8.60\mathrm{e}{-02}$ $1.76\mathrm{e}{-02}$
IVP $1.05\mathrm{e}{-02}$ $1.57\mathrm{e}{-02}$ $1.83\mathrm{e}{-05}$ $6.01\mathrm{e}{-05}$ $1.53\mathrm{e}{-02}$ $3.29\mathrm{e}{-02}$ $5.13\mathrm{e}{-03}$ $6.33\mathrm{e}{-03}$
Table 11: P-values for interpolation and extrapolation on Mouse Organogenesis from Table 3.
Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
$5.70\mathrm{e}{-03}$ $1.06\mathrm{e}{-03}$ $2.63\mathrm{e}{-02}$ $2.43\mathrm{e}{-01}$ $1.96\mathrm{e}{-03}$ $1.30\mathrm{e}{-03}$
@@ I.1 Ablation on $\lambda$
First, we conduct a sensitivity analysis on all three datasets for the parameter $\lambda$, which controls the relative importance of the two priors. All the other hyperparameters are kept constant. From Tables 12-16 and Figures 12-16, we observe that the best performance is usually achieved towards the extremities, at $\lambda=0$ or $\lambda=1$, with values near the latter dominating more often. We hypothesize that the spatial smoothness prior serves as a proxy for the spatial distance between cells from different slices and always carries relevant information, encoding the structural information present in the data. On the other hand, the informativeness of cell–cell communication patterns depends on how distinct the ligand–receptor features are at a given time step compared with those of neighboring ones. When these features remain highly similar across consecutive timesteps, they contribute little to the discriminative signal that the OT objective can leverage. Consequently, the influence of communication priors is strongly dataset- and timestep-dependent. This effect is also reflected in the observation that settings with $\lambda=0$ tend to perform worse when using Next-Step sampling—where local, immediate effects dominate—than under IVP sampling, which integrates information over the entire preceding trajectory. We therefore recommend experimenting with different values of $\lambda$ (e.g., 0, $0.8$, or $1$) depending on the specific use case and context. Due to ContextFlow’s scalability, hyperparameter exploration can be performed efficiently, enabling rapid assessment of the effect of $\lambda$.
@@ I.2 Ablation on $r$
We examine the effect of varying the neighborhood radius $r$ used to define the boundary for computing the spatial smoothness prior. We evaluate two settings for Brain Regeneration: $\lambda=1$ and $\lambda=0.8$, corresponding to using only the spatial prior and a setting with a modest contribution from the cell–cell communication prior. From Tables 17-20 and their corresponding Figures 17-20, we observe that the optimal neighborhood radius tends to lie toward the lower end of the tested range. Radii smaller than this optimum degrade performance by failing to capture sufficient local context, resulting in neighborhood means that are overly similar to individual cellular profiles. Conversely, increasing the radius beyond the optimal range also reduces performance, as the neighborhood begins to include cells from distinct types or spatial regions, thereby diluting the local signal. While certain deviations from this trend occur, likely reflecting underlying biological complexity, this behavior is consistent with the trade-off between spatial specificity and contextual coverage inherent to neighborhood-based priors. For our case, we set the radius by considering the timestep with the fewest cells, dividing it by $2$ (to account for different hemispheres), and dividing by the number of cell types present in that timestep. For the dataset considered in this study, Stage $44$ had the fewest number of cells, $1400$, with approximately $10$ cell types. We thus set the radius at $50$ in our studies.
@@ I.3 Ablation on $\epsilon$
Additionally, we conduct an ablation study on the Brain Regeneration dataset by varying the parameter $\epsilon$, which weights the entropic term in the OT objective. The ContextFlow configuration we consider here is CTF-H ($\lambda=1$), which includes only the spatial smoothness prior. As observed from Tables 21 and 22, as well as their corresponding Figures 21 and 22, drastically decreasing $\epsilon$ results in the OT formulation to ignore the relative entropic term containing the prior information and only to consider the transport cost, resulting in higher Wasserstein values. Furthermore, in accordance with results from Theorem 4.1, increasing $\epsilon$ too much still does not drastically degrade the performance, as the prior matrix $M$ acts as a soft filter and prohibits uniform couplings. While setting $\epsilon$, one must look at the Gibbs kernel used in the Sinkhorn Algorithm $\exp(-\mathbf{C}/\epsilon)$, since a small $\epsilon$ can cause potential numerical issues. We thus set $\epsilon$ by examining the order of the cost matrix $\mathbf{C}$, and, for the studies above, we set it to $100$ based on the median of all elements in the cost matrix.
@@
0 $1.508\pm 0.047$ $1.386\pm 0.088$ $6.986\pm 1.296$ $23.807\pm 7.516$
0.2 $1.614\pm 0.081$ $1.642\pm 0.136$ $6.907\pm 1.471$ $24.675\pm 9.126$
0.5 $1.638\pm 0.069$ $1.676\pm 0.114$ $7.749\pm 1.377$ $32.480\pm 8.322$
0.8 $1.617\pm 0.042$ $1.680\pm 0.094$ $7.645\pm 2.376$ $29.854\pm 10.015$
1 $1.906\pm 0.071$ $1.892\pm 0.092$ $6.378\pm 1.813$ $11.845\pm 6.567$
@@
Table 16: Interpolation for holdout timestep 2 with IVP Sampling on the Liver Regeneration dataset.
@@
Table 19: Extrapolation for the last holdout timestep with CTF-H at $\lambda=0.8$ on Brain Regeneration.
@@
Table 20: Interpolation for the middle holdout timestep with CTF-H at $\lambda=0.8$ on Brain Regeneration.